yallbabynene08:

Which expression is equivalent to 2 + 11w? Group of answer choices 2w x 11w 2w + 11w 2w + 11 11w + 2

snowflake0531:

The commutative property or whatever says a+b=b+a so 2+11w = ?

yallbabynene08:

um

yallbabynene08:

ohhhhhh

snowflake0531:

which one

yallbabynene08:

nvm